OpenStudy (anonymous):

[10.03] Which number equals the square root of 108.? 3 the square root of 3. 6 the square root of 2. 6 the square root of 3. 3 the square root of 6.

OpenStudy (anonymous):

\[\sqrt{108} = \sqrt{a}\sqrt{b^2} = b\sqrt{a}\]

OpenStudy (anonymous):

I am lost

OpenStudy (anonymous):

Hint: What are the factors of 108? More importantly, are there any factors which are square numbers?

OpenStudy (anonymous):

6 3 is the answer